calculate the emf of the following cell at 298K
H2+1/2O2=H2O  delta G0= -226.8kj/mol
H2O=H++OH-       delta G0= 76.2kj/mol

Grade:12th pass

2 Answers

Neeraj Joshi
9 years ago
let us take the frst equation as equation 1 and second as equation 2
Now,
add eq1+2 eq2
H2O + ½ O2 +H2 = 2H+ + 20H-
Then delta G will b 
-226.8+2(76.2)
-74.4kj/mol
as delta G=-nFE
E=-G/nF
E=74400/2*96500
E=0.385Jmol-1
ANIMESH
9 years ago
but the answer is 1.2 volts